OpenStudy (anonymous):

Solve for y: 3x + 2y < 12. Answer y < -3/2x + 6 y < -2/3x + 4 y > 3/2 x + 6 y < 3/2x + 4

OpenStudy (anonymous):

@MiaKitty

OpenStudy (anonymous):

I got C.

OpenStudy (anonymous):

Hmm...I know how to do this, but the thing is kinda of confusing me. I think it is the morning, I think C might be right. :)

OpenStudy (anonymous):

It alright (:

OpenStudy (anonymous):

See if this helps you... 3x+2y=12 2y=-3x+12 y=-3x+12/2 y=-3x/2+6